# The evolution and detection of vector superradiant instabilities

### v1

@@ II.4 Evolution equations
The superradiance results in the energy flux and the angular momentum flux through the BH horizon, which are given by,
@@
Superradiance of ${}^{3}l_{j,m}^{n}$ mode exists only when the BH spin is above this value. With the presence of the ${}^{3}l_{j,m}^{n}$ mode, there is an attractor of the BH spin. If BH spin is below this value, the angular momentum flows into the horizon. As a result, the BH spin increases while the ${}^{3}l_{j,m}^{n}$ mode shrinks. Conversely, if the BH spin is above this critical value, the angular momentum flows from the BH to the condensate, reducing the BH spin while increasing the size of the ${}^{3}l_{j,m}^{n}$ mode. In the limit of $M\mu\ll 1$, Eq. (38) has the expanded form,
@@
Meanwhile, the ${}^{3}l_{j,m}^{n}$ mode loses its energy and angular momentum continuously via GW emission. The two loss-rates are related by,
@@
Finally, the evolution equations of the BH-condensate system are,
@@
$\dot{J}_{\mathrm{v}}^{(nljm)}$ $=\dot{J}_{\mathrm{H}}^{(nljm)}-\dot{J}_{\mathrm{GW}}^{(nljm)}.$ (41d)
where $J$ and $M$ are the angular momentum and mass of the BH, respectively. And $J^{(nljm)}_{\mathrm{v}}$ is the angular momentum of the ${}^{3}l_{j,m}^{n}$ mode,
In obtaining these evolution equations, we have made the following assumptions,
  • The superradiance rate in Eq. (23) and the GW emission fluxes in Eqs. (26) and (28) requires $M\mu\ll 1$.
@@ III Evolution with a single dominant mode
The dominant modes have $l=j-1$. Then the superradiance rate in Eq. (23) is proportional to $(M\mu)^{4j+3}$, while the GW emission rate in Eq. (26) is proportional to $(M\mu)^{4j+6}$. In the limit of $M\mu\ll 1$, the superradiance process is much faster than the energy dissipation. This hierarchy results in two well-separated phases in the evolution of the dominant modes. In the first phase, the condensate extracts energy and angular momentum from the host BH. As a result, the BH spin drops until reaching the critical value $a_{*\text{c}}$ and the superradiance terminates. The evolution of this phase is governed by superradiance while GW emission can be safely ignored. We refer to this phase as spin-down phase. Then in the second phase, the condensate slowly loses mass and angular momentum via GW emission. The BH mass and spin are unchanged, which is an attractor mathematically. We refer to this phase as attractor phase.
@@ III.1 Analytic approximation
In the spin-down phase, the GW emission can be safely ignored. The initial BH mass and spin are $M_{0}$ and $a_{*0}$, respectively. The initial mass of the condensate is $M_{\mathrm{v0}}^{(0011)}$. Then the condensate extracts energy and angular momentum from the host BH. The spin-down phase ends at time $t_{1}$, when the BH spin decreases to $a^{(0011)}_{*\text{c}}$ in Eq. (38) and the BH mass drops to $M_{1}$. At the same time, the condensate reaches its maximum mass $M^{(0011)}_{\text{v,max}}$. The conservations of energy and angular momentum are,
@@
where $\omega_{0011}$ is calculated in Eq. (20) with $\alpha$ replaced by $M_{1}\mu$. By fitting the numerical results from solving the conservation equations, the $M_{\mathrm{v,max}}^{(0011)}$ can be expressed as a power series of $M_{0}\mu$,
@@
The condensate mass at the end of the spin-down phase can also be calculated by solving Eq. (41c) with $\dot{E}^{(nljm)}_{\text{GW}}$ setting to zero. The difficulty lies in the fact that $\dot{E}_{\text{H}}^{(0011)}$ depends on $a_{*}$ via the superradiance rate $\Gamma_{0011}$. From Fig. 3, $\Gamma_{0011}$ depends only weakly on $a_{*}$ except in the close neighbourhood of $a_{*\text{c}}$. Consequently, one could replace $a_{*}$ in $\Gamma_{0011}$ by $a_{*0}$ as a good approximation. The equation is then trivial to solve, resulting in exponential growth of the condensate mass,
$M_{\mathrm{v,max}}^{(0011)}\approx M_{\mathrm{v0}}^{(0011)}e^{2\Gamma_{0011}\tau_{\mathrm{sr}}^{(0011)}},$ (44)
where $\tau_{\text{sr}}$ is the duration of the spin-down phase. Then the duration of this phase can be calculated as,
$\begin{split}\frac{\tau_{\mathrm{sr}}^{(0011)}}{M_{0}}&\approx\frac{1}{2M_{0}\Gamma_{0011}}\log\frac{M_{\mathrm{v,max}}^{(0011)}}{M_{\mathrm{v0}}^{(0011)}}.\end{split}$ (45)
For the order of magnitude, one could simply set the logarithm to 10, which gives,
$\tau_{\mathrm{sr}}^{(0011)}\sim 20.5\ \mathrm{min}\left(\frac{M_{0}}{10M_{\odot}}\right)\left(\frac{0.1}{M_{0}\mu}\right)^{7}\left(\frac{1}{a_{*0}}\right).$ (46)
The attractor phase starts when the condensate reaches its maximum mass. At this moment, the BH spin is almost $a_{*\text{c}}^{(0011)}$ and superradiance is negligible. The BH mass and spin are then unchanged in this phase. Meanwhile, the condensate dissipates energy and angular momentum via GW emission. The condensate mass in this phase can be solved from Eq. (41c) with $\dot{E}_{\text{H}}^{(nljm)}$ set to zero. Inserting Eq. (26) and one arrives at,
@@
When $t\gg\tau_{\text{GW}}^{(0011)}$, since $\tau_{\text{GW}}^{(0011)}\gg\tau_{\text{sr}}^{(0011)}\sim t_{1}$, the expression reduces to,
@@ III.2.1 The effect of initial BH spin
Figure 5: The evolutions of the condensate mass (panel a), BH mass (panel b), and BH spin (panel c) as functions of time with different BH initial spins $a_{*0}$, where only ${}^{3}S_{1,1}^{0}$ mode is included in the evolution. The initial mass coupling is $M_{0}\mu=0.1$ and the initial mass of the condensate is $10^{-5}M_{0}$. The solid curves are the numerical solutions from solving the evolution equations in Eqs. (41). The dashed horizontal lines in the three panels are from the analytic approximations in Eqs. (43), (42a) and (39), respectively. For completeness, the evolution with initial BH spin $a_{*0}=0.25$ is also shown, which does not satisfy the superradiance condition $\omega_{0011}<\Omega_{\mathrm{H}}$.
@@
$M_{\mathrm{v},\max}^{(0011)}/M_{0}$ $M_{1}/M_{0}$ $a_{*\mathrm{c}}$ $\tau_{\mathrm{sr}}^{(0011)}/M_{0}$ $\tau_{\mathrm{GW}}^{(0011)}/M_{0}$ $M_{\mathrm{v},\max}^{(0011)}/M_{0}$ $M_{1}/M_{0}$ $a_{*\mathrm{c}}$ $\tau_{\mathrm{sr}}^{(0011)}/M_{0}$ $\tau_{\mathrm{GW}}^{(0011)}/M_{0}$
@@
Table 1: Comparison of the analytic approximations to the numerical solutions for the important quantities with different initial BH spins $a_{*0}$, where only ${}^{3}S_{1,1}^{0}$ mode is included in the evolution. The initial mass coupling is $M_{0}\mu=0.1$ and the initial mass of the condensate is $10^{-5}M_{0}$. The numerical solutions are from solving Eqs. (41). The estimates of the maximum condensate mass $M_{\text{v,max}}^{(0011)}$, the final BH mass $M_{1}$, the final BH spin $a_{*\text{c}}$, the superradiance time scale $\tau_{\mathrm{sr}}^{(0011)}$ and the GW emission time scale $\tau_{\mathrm{GW}}^{(0011)}$ are calculated with Eqs. (43), (42a), (39), (45) and (48), respectively.
@@
The analytic approximations are compared to the numerical values in table 1. The errors are all less than 10% except for $\tau_{\text{sr}}$. Nonetheless, Eq. (45) still gives the correct order of magnitude of $\tau_{\text{sr}}$.
@@ III.2.2 The effect of initial condensate mass values
Figure 6: The evolutions of the condensate mass (panel a), BH mass (panel b), and BH spin (panel c) as functions of time with different initial condensate masses $M_{\text{v0}}^{(0011)}$, where only ${}^{3}S_{1,1}^{0}$ mode is included in the evolution. The initial mass coupling is $M_{0}\mu=0.1$ and the initial BH spin is $a_{*0}=0.99$. The solid curves are the numerical solutions from solving the evolution equations in Eqs. (41). The dashed horizontal lines in the three panels are from the analytic approximations in Eqs. (43), (42a) and (39), respectively.
@@
$M_{\mathrm{v},\max}^{(0011)}/M_{0}$ $M_{1}/M_{0}$ $a_{*\mathrm{c}}$ $\tau_{\mathrm{sr}}^{(0011)}/M_{0}$ $\tau_{\mathrm{GW}}^{(0011)}/M_{0}$ $M_{\mathrm{v},\max}^{(0011)}/M_{0}$ $M_{1}/M_{0}$ $a_{*\mathrm{c}}$ $\tau_{\mathrm{sr}}^{(0011)}/M_{0}$ $\tau_{\mathrm{GW}}^{(0011)}/M_{0}$
@@
Table 2: Comparison of the analytic approximations to the numerical solutions for the important quantities with different initial condensate mass $M_{\text{v0}}^{(0011)}$, where only ${}^{3}S_{1,1}^{0}$ mode is included in the evolution. The initial mass coupling is $M_{0}\mu=0.1$ and the initial BH spin is $a_{*0}=0.99$. The numerical solutions are from solving Eqs. (41). The estimates of the maximum condensate mass $M_{\text{v,max}}^{(0011)}$, the final BH mass $M_{1}$, the final BH spin $a_{*\text{c}}$, the superradiance time scale $\tau_{\mathrm{sr}}^{(0011)}$ and the GW emission time scale $\tau_{\mathrm{GW}}^{(0011)}$ are calculated with Eqs. (43), (42a), (39), (45) and (48), respectively.
Fig. 6 shows the condensate mass, the BH mass and spin as functions of time. The fixed parameters are $a_{*0}=0.99$ and $M_{0}\mu=0.1$. Five different initial condensate masses $M_{\mathrm{v0}}^{(0011)}/M_{0}=10^{-1},10^{-2},10^{-3},10^{-4},10^{-5}$ are then compared. The comparison of the analytic approximations and the numerical results are further presented in table 2.
@@
In the attractor phase, the BH mass and spin stay unchanged, with values predicted by the analytic approximations. In particular, the BH mass and spin are independent of the initial condensate mass. In this phase, the condensate mass drops monotonically due to GW emission. When $t$ is larger than $\sim 10^{11}M_{0}$, the five curves in Fig. 6(a) merges, which is anticipated in Eq. (49).
@@ III.2.3 The effect of initial mass couplings
Figure 7: The evolutions of the condensate mass (panel a), BH mass (panel b), and BH spin (panel c) as functions of time with different initial mass couplings $M_{0}\mu$, where only ${}^{3}S_{1,1}^{0}$ mode is included in the evolution. The initial BH spin is $a_{*0}=0.99$ and the initial condensate mass is $M_{\mathrm{v0}}^{(0011)}=10^{-5}M_{0}$. The solid curves are the numerical solutions from solving the evolution equations in Eqs. (41). The dashed horizontal lines in the three panels are from the analytic approximations in Eqs. (43), (42a) and (39), respectively.
@@
$M_{\mathrm{v},\max}^{(0011)}/M_{0}$ $M_{1}/M_{0}$ $a_{*\mathrm{c}}$ $\tau_{\mathrm{sr}}^{(0011)}/M_{0}$ $\tau_{\mathrm{GW}}^{(0011)}/M_{0}$ $M_{\mathrm{v},\max}^{(0011)}/M_{0}$ $M_{1}/M_{0}$ $a_{*\mathrm{c}}$ $\tau_{\mathrm{sr}}^{(0011)}/M_{0}$ $\tau_{\mathrm{GW}}^{(0011)}/M_{0}$
@@
Table 3: Comparison of the analytic approximations to the numerical solutions for the important quantities with different initial mass coupling $M_{0}\mu$, where only ${}^{3}S_{1,1}^{0}$ mode is included in the evolution. The initial BH spin is $a_{*0}=0.99$ and the initial mass of the condensate is $10^{-5}M_{0}$. The numerical solutions are from solving Eqs. (41). The estimates of the maximum condensate mass $M_{\text{v,max}}^{(0011)}$, the final BH mass $M_{1}$, the final BH spin $a_{*\text{c}}$, the superradiance time scale $\tau_{\mathrm{sr}}^{(0011)}$ and the GW emission time scale $\tau_{\mathrm{GW}}^{(0011)}$ are calculated with Eqs. (43), (42a), (39), (45) and (48), respectively.
Fig. 7 shows the condensate mass, the BH mass and spin as functions of time. Fixed parameters are $a_{*0}=0.99$ and $M_{\text{v0}}/M_{0}=10^{-5}$. Five different initial mass couplings $M_{0}\mu=0.01,0.05,0.1,0.3,0.4$ are compared. The comparison of the analytic approximations and the numerical results are further presented in table 3.

### v2

@@ II.4 Evolution equations
The superradiance results in the energy flux and the angular momentum flux through the BH horizon, which are given by
@@
Superradiance of the ${}^{3}l_{j,m}^{n}$ mode exists only when the BH spin is above this value. With the presence of the ${}^{3}l_{j,m}^{n}$ mode, there is an attractor of the BH spin. If BH spin is below this value, the angular momentum flows into the horizon. As a result, the BH spin increases while the ${}^{3}l_{j,m}^{n}$ mode shrinks. Conversely, if the BH spin is above this critical value, the angular momentum flows from the BH to the condensate, reducing the BH spin while increasing the size of the ${}^{3}l_{j,m}^{n}$ mode. In the limit of $M\mu\ll 1$, Eq. (38) has the expanded form
@@
Meanwhile, the ${}^{3}l_{j,m}^{n}$ mode loses its energy and angular momentum continuously via GW emission. The two loss-rates are related by
@@
Finally, the evolution equations of the BH-condensate system are
@@
$\dot{J}_{\mathrm{v}}^{(nljm)}$ $=\dot{J}_{\mathrm{H}}^{(nljm)}-\dot{J}_{\mathrm{GW}}^{(nljm)},$ (41d)
where $J$ and $M$ are the angular momentum and mass of the BH, respectively, and $J^{(nljm)}_{\mathrm{v}}$ is the angular momentum of the ${}^{3}l_{j,m}^{n}$ mode.
In obtaining these evolution equations, we have made the following assumptions:
  • The superradiance rate in Eq. (23) and the GW emission fluxes in Eqs. (26) and (28) require $M\mu\ll 1$.
@@ III Evolution with a single dominant mode
The dominant modes have $l=j-1$. Then the superradiance rate in Eq. (23) is proportional to $(M\mu)^{4j+3}$, while the GW emission rate in Eq. (26) is proportional to $(M\mu)^{4j+6}$. In the limit of $M\mu\ll 1$, the superradiance process is much faster than the energy dissipation. This hierarchy results in two well-separated phases in the evolution of the dominant modes. In the first phase, the condensate extracts energy and angular momentum from the host BH. As a result, the BH spin drops until reaching the critical value $a_{*\text{c}}$ and the superradiance terminates. The evolution of this phase is governed by superradiance while GW emission can be safely ignored. We refer to this phase as the spin-down phase. Then in the second phase, the condensate slowly loses mass and angular momentum via GW emission. The BH mass and spin are unchanged, which is an attractor mathematically. We refer to this phase as the attractor phase.
@@ III.1 Analytic approximation
In the spin-down phase, the GW emission can be safely ignored. The initial BH mass and spin are $M_{0}$ and $a_{*0}$, respectively. The initial mass of the condensate is $M_{\mathrm{v0}}^{(0011)}$. Then the condensate extracts energy and angular momentum from the host BH. The spin-down phase ends at time $t_{1}$, when the BH spin decreases to $a^{(0011)}_{*\text{c}}$ in Eq. (38) and the BH mass drops to $M_{1}$. At the same time, the condensate reaches its maximum mass $M^{(0011)}_{\text{v,max}}$. The conservations of energy and angular momentum are
@@
where $\omega_{0011}$ is calculated in Eq. (20) with $\alpha$ replaced by $M_{1}\mu$. By fitting the numerical results from solving the conservation equations, $M_{\mathrm{v,max}}^{(0011)}$ can be expressed as a power series of $M_{0}\mu$
@@
The condensate mass at the end of the spin-down phase can also be calculated by solving Eq. (41c) with $\dot{E}^{(nljm)}_{\text{GW}}$ set to zero. The difficulty lies in the fact that $\dot{E}_{\text{H}}^{(0011)}$ depends on $a_{*}$ via the superradiance rate $\Gamma_{0011}$. From Fig. 3, $\Gamma_{0011}$ depends only weakly on $a_{*}$ except in the close neighborhood of $a_{*\text{c}}$. Consequently, one could replace $a_{*}$ in $\Gamma_{0011}$ by $a_{*0}$ as a good approximation. The equation is then trivial to solve, resulting in exponential growth of the condensate mass,
$M_{\mathrm{v,max}}^{(0011)}\approx M_{\mathrm{v0}}^{(0011)}e^{2\Gamma_{0011}t_{1}}.$ (44)
Then the duration of this phase can be calculated as
$t_{1}\approx\tau_{\mathrm{sr}}^{(0011)}\log\frac{M_{\mathrm{v,max}}^{(0011)}}{M_{\mathrm{v0}}^{(0011)}},$ (45)
where $\tau_{\text{sr}}^{(0011)}$ is the superradiant timescale,
$\begin{split}\tau_{\mathrm{sr}}^{(0011)}&=\frac{1}{2\Gamma_{0011}},\\ &\approx 2.05\ \mathrm{min}\left(\frac{M_{0}}{10M_{\odot}}\right)\left(\frac{0.1}{M_{0}\mu}\right)^{7}\left(\frac{1}{a_{*0}}\right).\end{split}$ (46)
The attractor phase starts when the condensate reaches its maximum mass. At this moment, the BH spin is almost $a_{*\text{c}}^{(0011)}$ and superradiance is negligible. The BH mass and spin are then unchanged in this phase. Meanwhile, the condensate dissipates energy and angular momentum via GW emission. The condensate mass in this phase can be solved from Eq. (41c) with $\dot{E}_{\text{H}}^{(nljm)}$ set to zero. Inserting Eq. (26), one arrives at
@@
When $t\gg\tau_{\text{GW}}^{(0011)}$, since $\tau_{\text{GW}}^{(0011)}\gg\tau_{\text{sr}}^{(0011)}\sim t_{1}$, the expression reduces to
@@ III.2.1 The effect of initial BH spin
Figure 5: The evolutions of the condensate mass (panel a), BH mass (panel b), and BH spin (panel c) as functions of time with different BH initial spins $a_{*0}$, where only the ${}^{3}S_{1,1}^{0}$ mode is included in the evolution. The initial mass coupling is $M_{0}\mu=0.1$ and the initial mass of the condensate is $10^{-5}M_{0}$. The solid curves are the numerical solutions from solving the evolution equations in Eqs. (41). The dashed horizontal lines in the three panels are from the analytic approximations in Eqs. (43), (42a) and (39), respectively. For completeness, the evolution with initial BH spin $a_{*0}=0.25$ is also shown, which does not satisfy the superradiance condition $\omega_{0011}<\Omega_{\mathrm{H}}$.
@@
$M_{\mathrm{v},\max}^{(0011)}/M_{0}$ $M_{1}/M_{0}$ $a_{*\mathrm{c}}$ $t_{1}/M_{0}$ $\tau_{\mathrm{GW}}^{(0011)}/M_{0}$ $M_{\mathrm{v},\max}^{(0011)}/M_{0}$ $M_{1}/M_{0}$ $a_{*\mathrm{c}}$ $t_{1}/M_{0}$ $\tau_{\mathrm{GW}}^{(0011)}/M_{0}$
@@
Table 1: Comparison of the analytic approximations to the numerical solutions for the important quantities with different initial BH spins $a_{*0}$, where the only the ${}^{3}S_{1,1}^{0}$ mode is included in the evolution. The initial mass coupling is $M_{0}\mu=0.1$ and the initial mass of the condensate is $10^{-5}M_{0}$. The numerical solutions are from solving Eqs. (41). The estimates of the maximum condensate mass $M_{\text{v,max}}^{(0011)}$, the final BH mass $M_{1}$, the final BH spin $a_{*\text{c}}$, the duration of the spin-down phase $t_{1}$ and the GW emission time scale $\tau_{\mathrm{GW}}^{(0011)}$ are calculated with Eqs. (43), (42a), (39), (45) and (48), respectively.
@@
The analytic approximations are compared to the numerical values in Table 1. The errors are all less than 10% except for $t_{1}$. Nonetheless, Eq. (45) still gives the correct order of magnitude of $t_{1}$.
@@ III.2.2 The effect of initial condensate mass values
Figure 6: The evolutions of the condensate mass (panel a), BH mass (panel b), and BH spin (panel c) as functions of time with different initial condensate masses $M_{\text{v0}}^{(0011)}$, where only the ${}^{3}S_{1,1}^{0}$ mode is included in the evolution. The initial mass coupling is $M_{0}\mu=0.1$ and the initial BH spin is $a_{*0}=0.99$. The solid curves are the numerical solutions from solving the evolution equations in Eqs. (41). The dashed horizontal lines in the three panels are from the analytic approximations in Eqs. (43), (42a) and (39), respectively.
@@
$M_{\mathrm{v},\max}^{(0011)}/M_{0}$ $M_{1}/M_{0}$ $a_{*\mathrm{c}}$ $t_{1}/M_{0}$ $\tau_{\mathrm{GW}}^{(0011)}/M_{0}$ $M_{\mathrm{v},\max}^{(0011)}/M_{0}$ $M_{1}/M_{0}$ $a_{*\mathrm{c}}$ $t_{1}/M_{0}$ $\tau_{\mathrm{GW}}^{(0011)}/M_{0}$
@@
Table 2: Comparison of the analytic approximations to the numerical solutions for the important quantities with different initial condensate mass $M_{\text{v0}}^{(0011)}$, where only the ${}^{3}S_{1,1}^{0}$ mode is included in the evolution. The initial mass coupling is $M_{0}\mu=0.1$ and the initial BH spin is $a_{*0}=0.99$. The numerical solutions are from solving Eqs. (41). The estimates of the maximum condensate mass $M_{\text{v,max}}^{(0011)}$, the final BH mass $M_{1}$, the final BH spin $a_{*\text{c}}$, the duration of the spin-down phase $t_{1}$ and the GW emission time scale $\tau_{\mathrm{GW}}^{(0011)}$ are calculated with Eqs. (43), (42a), (39), (45) and (48), respectively.
Figure 6 shows the condensate mass, the BH mass and spin as functions of time. The fixed parameters are $a_{*0}=0.99$ and $M_{0}\mu=0.1$. Five different initial condensate masses $M_{\mathrm{v0}}^{(0011)}/M_{0}=10^{-1},10^{-2},10^{-3},10^{-4},10^{-5}$ are then compared. The comparison of the analytic approximations and the numerical results are further presented in Table 2.
@@
In the attractor phase, the BH mass and spin stay unchanged, with values predicted by the analytic approximations. In particular, the BH mass and spin are independent of the initial condensate mass. In this phase, the condensate mass drops monotonically due to GW emission. When $t$ is larger than $\sim 10^{11}M_{0}$, the five curves in Fig. 6(a) merge, which is anticipated in Eq. (49).
@@ III.2.3 The effect of initial mass couplings
Figure 7: The evolutions of the condensate mass (panel a), BH mass (panel b), and BH spin (panel c) as functions of time with different initial mass couplings $M_{0}\mu$, where only the ${}^{3}S_{1,1}^{0}$ mode is included in the evolution. The initial BH spin is $a_{*0}=0.99$ and the initial condensate mass is $M_{\mathrm{v0}}^{(0011)}=10^{-5}M_{0}$. The solid curves are the numerical solutions from solving the evolution equations in Eqs. (41). The dashed horizontal lines in the three panels are from the analytic approximations in Eqs. (43), (42a) and (39), respectively.
@@
$M_{\mathrm{v},\max}^{(0011)}/M_{0}$ $M_{1}/M_{0}$ $a_{*\mathrm{c}}$ $t_{1}/M_{0}$ $\tau_{\mathrm{GW}}^{(0011)}/M_{0}$ $M_{\mathrm{v},\max}^{(0011)}/M_{0}$ $M_{1}/M_{0}$ $a_{*\mathrm{c}}$ $t_{1}/M_{0}$ $\tau_{\mathrm{GW}}^{(0011)}/M_{0}$
@@
Table 3: Comparison of the analytic approximations to the numerical solutions for the important quantities with different initial mass coupling $M_{0}\mu$, where only the ${}^{3}S_{1,1}^{0}$ mode is included in the evolution. The initial BH spin is $a_{*0}=0.99$ and the initial mass of the condensate is $10^{-5}M_{0}$. The numerical solutions are from solving Eqs. (41). The estimates of the maximum condensate mass $M_{\text{v,max}}^{(0011)}$, the final BH mass $M_{1}$, the final BH spin $a_{*\text{c}}$, the duration of the spin-down phase $t_{1}$ and the GW emission time scale $\tau_{\mathrm{GW}}^{(0011)}$ are calculated with Eqs. (43), (42a), (39), (45) and (48), respectively.
Figure 7 shows the condensate mass, the BH mass and spin as functions of time. Fixed parameters are $a_{*0}=0.99$ and $M_{\text{v0}}/M_{0}=10^{-5}$. Five different initial mass couplings $M_{0}\mu=0.01,0.05,0.1,0.3,0.4$ are compared. The comparison of the analytic approximations and the numerical results are further presented in Table 3.